# ARC-AGI-2: A New Challenge for Frontier AI Reasoning Systems

### v1

@@ 1 ARC-AGI-1: 2019-2024 history
The Abstraction and Reasoning Corpus (ARC), later referred to as ARC-AGI-1, was introduced by François Chollet in the 2019 paper “On the Measure of Intelligence”  [10]. It represented a significant departure from traditional AI benchmarks, which often focused on specific skills or knowledge recall within large datasets. Instead, ARC-AGI-1 was designed to evaluate a more general, human-like form of fluid intelligence — the ability to reason and solve novel problems efficiently, independent of extensive prior experience or domain-specific training.
@@ 1.1 The dataset
The original public dataset contained 400 training tasks and 400 evaluation tasks. A Private Evaluation set of 100 tasks was reserved for competition leaderboards, and a further 100-task "Semi-Private" Evaluation set was introduced in mid-2024 for verifying closed-source models  [1].
@@
    It minimizes the need for prior knowledge. ARC-AGI tasks only require a core set of innate human cognitive priors, such as object persistence, goal-directedness, elementary number sense (counting), and basic geometric and topological concepts (connectivity, symmetry)  [10]. Crucially, it requires absolutely no specialized world knowledge (e.g., historical facts) nor language to solve, making it distinct from knowledge-intensive benchmarks commonly used for evaluating LLMs.
@@ 1.2 Prior competitions
    2020: First ARC-AGI Kaggle competition (\$20,000 in prizes)  [3]
@@
    2024: ARC Prize 2024 (\$1,100,000 in prizes)  [7]
For several years following its introduction, progress remained slow. The winning entry in the 2020 competition achieved a score of 20% on the Private Evaluation set using program synthesis techniques  [3]. Despite the rapid scaling of large language models (LLMs) between 2020 and early 2024 (during which base LLMs were scaled up by a factor of more than 10,000x), the state-of-the-art ARC-AGI scores hovered around 34%  [7], well below estimated human performance.
The landscape shifted rapidly during the ARC Prize 2024 competition, the first competition run by ARC Prize Foundation  [9]. The highest score achieved during the competition on the Private Evaluation set reached 55.5% (by the MindsAI team), although this solution was not open-sourced and was thus ineligible for prizes. The winning team with an eligible submission, “the ARChitects”, achieved 53.5%  [7]. This significant jump from the previous state-of-the-art highlighted the effectiveness of novel approaches spurred by the competition, in particular Test-Time Adaptation (TTA) approaches, in which a model dynamically adapts at test time via either test-time search (e.g., chain-of-thought synthesis or symbolic program synthesis) or test-time gradient descent. These approaches are discussed in the ARC Prize 2024 Technical Report  [8].
Shortly following the conclusion of the competition in late 2024, a preview version of OpenAI’s “o3” model, in private testing, demonstrated remarkable performance on the ARC-AGI-1 Semi-Private Evaluation dataset. Using significant computational resources at test time, this model achieved scores of 76% (low-compute; estimated cost: \$200 per task) and 88% (high-compute; estimated cost: \$20,000 per task)  [12], surpassing the nominal human baseline for the first time. Subsequent publicly released versions of o3 showed lower, albeit still strong, performance (e.g., 53% for o3-medium on the Semi-Private set)  [11], the preview result indicated that ARC-AGI-1, while still challenging, was approaching saturation under the test-time adaptation paradigm using very large amounts of compute power.
@@ 2 Limitations of ARC-AGI-1
Previous competitions have served to highlight several key limitations  [8] of ARC-AGI-1:
@@
While the original ARC-AGI private tasks were confirmed solvable by humans (initial testing by two individuals yielded scores of 97% and 98%), and while some third-party studies have investigated the performance of Mechanical Turkers on the public datasets (such as NYU’s "Fast and flexible: Human program induction in abstract reasoning tasks"  [13] and "H-ARC: A Robust Estimate of Human Performance on the Abstraction and Reasoning Corpus Benchmark"  [14]), there was no official first-party human baseline score derived under consistent conditions for the hidden tasks used in competitions.
@@ 5 Task selection process
After these selection and validation processes, final ARC-AGI-2 test pairs were solved, on average, by 75% of people who attempted them. The average test-taker solved 66% of tasks they attempted. 100% of ARC-AGI-2 tasks were solved by at least two people (many were solved by more) in two attempts or less.
@@ 6 State-of-the-art
Baseline model performance on ARC-AGI-2 was generated with the publicly available Model Baseline repository  [6]. Models were evaluated on the Semi-Private Evaluation set.
Complete updated scores are hosted on the official ARC-AGI Leaderboard  [1].
@@ 8.1 Kaggle competition
    Grand Prize (\$700,000): Awarded to the first team(s) to get $\geq$85accuracy on the hidden ARC-AGI-2 Private Evaluation set.
@@ 8.2 Public Leaderboard
The public ARC-AGI Leaderboard  [1] offers a snapshot of how frontier closed-weight models perform on the benchmark. While these submissions do not qualify for ARC Prize 2025 awards, they demonstrate what publicly accessible systems can currently achieve.
@@
ARC Prize also restricts which companies and models are eligible for testing. The public testing policy is provided on the ARC Prize website at arcprize.org/policy [2].

### v2

@@ 1 ARC-AGI-1: 2019-2024 history
The Abstraction and Reasoning Corpus (ARC), later referred to as ARC-AGI-1, was introduced by François Chollet in the 2019 paper “On the Measure of Intelligence” [chollet2019measure]. It represented a significant departure from traditional AI benchmarks, which often focused on specific skills or knowledge recall within large datasets. Instead, ARC-AGI-1 was designed to evaluate a more general, human-like form of fluid intelligence — the ability to reason and solve novel problems efficiently, independent of extensive prior experience or domain-specific training.
@@ 1.1 The dataset
The original public dataset contained 400 training tasks and 400 evaluation tasks. A Private Evaluation set of 100 tasks was reserved for competition leaderboards, and a further 100-task "Semi-Private" Evaluation set was introduced in mid-2024 for verifying closed-source models [arcprize:leaderboard].
@@
    It minimizes the need for prior knowledge. ARC-AGI tasks only require a core set of innate human cognitive priors, such as object persistence, goal-directedness, elementary number sense (counting), and basic geometric and topological concepts (connectivity, symmetry) [chollet2019measure]. Crucially, it requires absolutely no specialized world knowledge (e.g., historical facts) nor language to solve, making it distinct from knowledge-intensive benchmarks commonly used for evaluating LLMs.
@@ 1.2 Prior competitions
    2020: First ARC-AGI Kaggle competition (\$20,000 in prizes)  [kaggle2020]
@@
    2024: ARC Prize 2024 (\$1,100,000 in prizes)  [arcprize2024]
For several years following its introduction, progress remained slow. The winning entry in the 2020 competition achieved a score of 20% on the Private Evaluation set using program synthesis techniques  [kaggle2020]. Despite the rapid scaling of large language models (LLMs) between 2020 and early 2024 (during which base LLMs were scaled up by a factor of more than 10,000x), the state-of-the-art ARC-AGI scores hovered around 34%  [arcprize2024], well below estimated human performance.
The landscape shifted rapidly during the ARC Prize 2024 competition, the first competition run by ARC Prize Foundation [arcprizefoundation]. The highest score achieved during the competition on the Private Evaluation set reached 55.5% (by the MindsAI team), although this solution was not open-sourced and was thus ineligible for prizes. The winning team with an eligible submission, “the ARChitects”, achieved 53.5%  [arcprize2024]. This significant jump from the previous state-of-the-art highlighted the effectiveness of novel approaches spurred by the competition, in particular Test-Time Adaptation (TTA) approaches, in which a model dynamically adapts at test time via either test-time search (e.g., chain-of-thought synthesis or symbolic program synthesis) or test-time gradient descent. These approaches are discussed in the ARC Prize 2024 Technical Report  [arcprize2024technicalreport].
Shortly following the conclusion of the competition in late 2024, a preview version of OpenAI’s “o3” model, in private testing, demonstrated remarkable performance on the ARC-AGI-1 Semi-Private Evaluation dataset. Using significant computational resources at test time, this model achieved scores of 76% (low-compute; estimated cost: \$200 per task) and 88% (high-compute; estimated cost: \$20,000 per task)  [o3arcagi], surpassing the nominal human baseline for the first time. Subsequent publicly released versions of o3 showed lower, albeit still strong, performance (e.g., 53% for o3-medium on the Semi-Private set)  [o3analysis], the preview result indicated that ARC-AGI-1, while still challenging, was approaching saturation under the test-time adaptation paradigm using very large amounts of compute power.
@@ 2 Limitations of ARC-AGI-1
Previous competitions have served to highlight several key limitations  [arcprize2024technicalreport] of ARC-AGI-1:
@@
While the original ARC-AGI private tasks were confirmed solvable by humans (initial testing by two individuals yielded scores of 97% and 98%), and while some third-party studies have investigated the performance of Mechanical Turkers on the public datasets (such as NYU’s "Fast and flexible: Human program induction in abstract reasoning tasks"  [nyu2021] and "H-ARC: A Robust Estimate of Human Performance on the Abstraction and Reasoning Corpus Benchmark"  [nyu2024]), there was no official first-party human baseline score derived under consistent conditions for the hidden tasks used in competitions.
@@ 5 Task selection process
After selection and validation, 100% of ARC-AGI-2 tasks were solved by at least two independent, non-expert human testers drawn from the general public, with each task attempted by between 2 and 10 humans. Aggregating results by task, 75% of human attempts were successfully completed. Aggregated across all human evaluations and attempts, 66% of attempted test pairs were successfully completed. Together, these results establish that all tasks in ARC-AGI-2 are human-solvable.
@@ 6 State-of-the-art
Baseline model performance on ARC-AGI-2 was generated with the publicly available Model Baseline repository [arcprize:model_baseline]. Models were evaluated on the Semi-Private Evaluation set.
Complete updated scores are hosted on the official ARC-AGI Leaderboard [arcprize:leaderboard].
@@ 8.1 Kaggle competition
    Grand Prize (\$700,000): Awarded to the first team(s) to get $\geq$85% accuracy on the hidden ARC-AGI-2 Private Evaluation set.
@@ 8.2 Public Leaderboard
The public ARC-AGI Leaderboard [arcprize:leaderboard] offers a snapshot of how frontier closed-weight models perform on the benchmark. While these submissions do not qualify for ARC Prize 2025 awards, they demonstrate what publicly accessible systems can currently achieve.
@@
ARC Prize also restricts which companies and models are eligible for testing. The public testing policy is provided on the ARC Prize website at arcprize.org/policy [arcprize:policy].